### v2

# What’s Inside Your Diffusion Model? A Score-Based Riemannian Metric to Explore the Data Manifold

Abstract
Recent advances in diffusion models have demonstrated their remarkable ability to capture complex image distributions, but the geometric properties of the learned data manifold remain poorly understood. We address this gap by introducing a score-based Riemannian metric that leverages the Stein score function from diffusion models to characterize the intrinsic geometry of the data manifold without requiring explicit parameterization. Our approach defines a metric tensor in the ambient space that stretches distances perpendicular to the manifold while preserving them along tangential directions, effectively creating a geometry where geodesics naturally follow the manifold’s contours. We develop efficient algorithms for computing these geodesics and demonstrate their utility for both interpolation between data points and extrapolation beyond the observed data distribution. Through experiments on synthetic data with known geometry, Rotated MNIST, and complex natural images via Stable Diffusion, we show that our score-based geodesics capture meaningful transformations that respect the underlying data distribution. Our method consistently outperforms baseline approaches on perceptual metrics (LPIPS) and distribution-level metrics (FID, KID), producing smoother, more realistic image transitions. These results reveal the implicit geometric structure learned by diffusion models and provide a principled way to navigate the manifold of natural images through the lens of Riemannian geometry.
1 Introduction
The geometry of natural images can be studied by viewing them as points on a manifold in high-dimensional ambient – pixel – space. This perspective, commonly known as the manifold hypothesis [1, 2], suggests that despite the vast dimensionality of pixel space, natural images concentrate near a much lower-dimensional structure. Understanding this geometric structure is crucial for numerous applications including image synthesis, manipulation, and analysis [3, 4], visual perception [5, 6], and representation learning [7, 8, 9].
In this work, we propose a data-dependent metric derived from the Stein score function, which can be obtained from a diffusion model trained on a given set of images. Recent advances in diffusion models have demonstrated remarkable capabilities in capturing complex image distributions [10, 11], and several approaches have explored data-dependent metrics to uncover the geometrical properties of image manifolds [12, 13, 14]. However, these previous methods often require explicit manifold parameterization, are computationally intractable for high-dimensional data, or fail to capture the intrinsic geometric structure that reflects perceptual relationships in the data.
Our approach defines a Riemannian metric in the ambient – pixel – space, leveraging score functions from diffusion models to create a geometry that stretches perpendicular to the data manifold while preserving distances along it. This metric allows us to compute distances and geodesic paths between images directly in pixel space, capturing meaningful relationships that respect the underlying data distribution. Unlike the Euclidean distance, which treats all directions in pixel space equally, our score-induced metric accounts for the anisotropic nature of the image manifold, heavily penalizing movement in directions orthogonal to the manifold while preserving natural movement along tangential directions. By generating images along these geodesics, by exploiting a diffusion model, we can visualize and validate these transformations, providing unique insights into the structure of the learned image manifold.
The main contributions of our work are:
  • A novel score-based Riemannian metric that captures the geometry of the data manifold without requiring explicit parameterization
  • Efficient algorithms for computing geodesics and manifold extrapolation in high-dimensional image space
  • Empirical validation on both synthetic data with known geometry and complex image data
2 Background
2.1 The Manifold Hypothesis for Natural Images
The manifold hypothesis posits that high-dimensional data, such as natural images, approximately lies on a low-dimensional manifold embedded in the ambient space $\mathbb{R}^{N}$ [1, 2]. In this framework, images are represented as points in a $N$-dimensional pixel space, where each axis represents a pixel value. Despite the vast dimensionality of this space – often millions of dimensions for modern high-resolution images –, natural images occupy only a tiny fraction of it due to strong correlations and constraints that restrict the set of plausible pixel configurations [15].
Each set of images can be characterized by its own probability density function $p(\boldsymbol{x})$ defined over the pixel space $\mathbb{R}^{N}$. The key insight is that $p(\boldsymbol{x})$ concentrates its mass on a much lower-dimensional structure than the ambient dimension $N$, referred to as the support of the data distribution. Under the manifold hypothesis, this support approximates a lower-dimensional manifold $\mathcal{M}\subset\mathbb{R}^{N}$ (Figure 1 A). To capture the geometry of this implicit manifold, we employ data-induced metrics - Riemannian metrics defined in the ambient space that adapt locally according to the underlying probability density [16]. Our approach explores how ambient space deformation enables geodesics that naturally follow the manifold’s contours (Figure 1 B,C).
Figure 1: (A) Data points $\mathbf{x}$ sampled from a probability distribution $p(\mathbf{x})$ (left) concentrate on a lower-dimensional manifold $\mathcal{M}$ (right). (B) Linear interpolation (black dashed) versus geodesic interpolation (red curve) between MNIST digits. Geodesics follow the manifold surface, producing valid digit transitions, while linear paths yield superpositions. (C) Geometric deformation: data manifold $\mathcal{M}$ embedded in $\mathbb{R}^{N}$ (left) transforms flat Euclidean space into curved metric space ~$\tilde{\mathbb{R}}^{N}$ (middle, right). In this deformed space, geodesics of $\mathcal{M}$ can be computed directly as geodesics of $\tilde{\mathbb{R}^{N}}$.
2.2 Diffusion Models & Stein Score
Diffusion models have emerged as powerful generative approaches that gradually transform noise into structured data through an iterative denoising process [17, 10]. These models define a forward process $q$ that progressively adds Gaussian noise to data points $\mathbf{x}_{0}\sim p(\mathbf{x})$ according to a predefined schedule, creating a sequence of increasingly noisy versions $\mathbf{x}_{t}$ with $t$ representing the diffusion step:
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$ (1)
where $\bar{\alpha}_{t}$ represents cumulative noise schedule parameters. This process transforms any complex data distribution into a simple Gaussian distribution at the limit $t\rightarrow T$. The conditional distribution can be equivalently expressed in a sampling form, which makes the noise component explicit:
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (2)
where $\boldsymbol{\epsilon}$ is the standard Gaussian noise added during the forward diffusion process. This formulation highlights that at any timestep $t$, the noisy sample $\mathbf{x}_{t}$ is a combination of the original data point $\mathbf{x}_{0}$ and noise $\boldsymbol{\epsilon}$, with their proportions determined by the noise schedule.
Central to diffusion models is the score function $\mathbf{s}(\mathbf{x})=\nabla_{\mathbf{x}}\log p(\mathbf{x})$, which represents the gradient of the log probability density. For data concentrated on a low-dimensional manifold $\mathcal{M}\subset\mathbb{R}^{N}$, the score function has a crucial geometric interpretation: it points approximately normal to the data manifold, with magnitude increasing with distance from the manifold [13, 18]. When training diffusion models, the neural network is typically trained to predict the noise $\boldsymbol{\epsilon}$ that was added during the forward process. This prediction, denoted as $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$, aims to approximate the true noise $\boldsymbol{\epsilon}$ used to generate $\mathbf{x}_{t}$ from $\mathbf{x}_{0}$. Importantly, this noise prediction directly relates to the score of the noisy distribution through:
$\mathbf{s}_{t}(\mathbf{x}_{t})=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})=-\frac{\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)}{\sigma_{t}}$ (3)
where $\sigma_{t}=\sqrt{1-\bar{\alpha}_{t}}$ is the standard deviation of the noise at timestep $t$. This relationship, derived from the Tweedie-Robbins-Miyasawa formula [19, 20], enables us to extract geometric information about the underlying data manifold without requiring explicit parameterization.
3 Methods
3.1 Capturing Manifold Geometry via Ambient Metric Deformation: the Stein Score Metric Tensor
"Space tells matter how to move; matter tells spacetime how to curve" C.W. Misner et al. [21]
We define a data-induced metric through geometric deformation [22, 23], drawing inspiration from general relativity, where mass curves spacetime to create gravitational paths. Analogously, our approach deforms the ambient Euclidean space such that geodesics naturally adhere to the data manifold without requiring explicit parameterization (Figure 1 C). Rather than directly characterizing the manifold $\mathcal{M}$, we instead deform the geometry of $\mathbb{R}^{N}$ to accommodate the manifold’s presence—effectively creating a "gravitational pull" toward the data distribution.
To formalize this intuition, we leverage score vectors (normal to the manifold) to indirectly characterize the geometry.
[Definition] The Stein score metric tensor is a smooth map $g:\mathbb{R}^{N}\to\text{SPD}(N)$ (the space of symmetric positive-definite $N\times N$ matrices) defined at a point $\boldsymbol{x}\in\mathbb{R}^{N}$ as:
$g(\boldsymbol{x})=\mathbf{I}+\lambda\cdot\mathbf{s}(\boldsymbol{x})\mathbf{s}(\boldsymbol{x})^{T},$ (4)
where $\mathbf{I}$ is the $N\times N$ identity matrix, $\mathbf{s}(\boldsymbol{x})=\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x})$ is the score function, and $\lambda$ is a positive penalty parameter that controls the strength of penalization along normal directions.
The Stein score metric $g(\boldsymbol{x})$ equips the ambient space with a Riemannian structure respecting the underlying data manifold geometry. The inner product between vectors $\boldsymbol{u},\boldsymbol{v}\in\mathbb{R}^{N}$ at point $\boldsymbol{x}$ becomes: $<\boldsymbol{u},\boldsymbol{v}>_{g(\boldsymbol{x})}=\boldsymbol{u}^{T}\boldsymbol{v}+\lambda(\mathbf{s}(\boldsymbol{x})^{T}\boldsymbol{u})(\mathbf{s}(\boldsymbol{x})^{T}\boldsymbol{v})$, combining the standard Euclidean inner product with a term that penalizes movement normal to the manifold.
This construction has several desirable properties:
  1. 1.
    When $\lambda=0$, we recover the standard Euclidean metric;
  2. 2.
    As $\lambda$ increases, the metric becomes increasingly stretched in the direction of the score vector;
  3. 3.
    The metric remains positive definite for all values of $\lambda$ (see Appendix A).
3.2 Geometric computation on the data manifold
We now develop the mathematical foundation for understanding distances and shortest paths in our deformed ambient space.
Length of a curve and Energy functionals
A curve in the ambient space $\mathbb{R}^{N}$ is a smooth function $\gamma:[0,1]\rightarrow\mathbb{R}^{N}$. The length of this curve under our score-based metric is computed as:
[…]
Where $\dot{\gamma}(\tau)=\frac{d}{d\tau}{\gamma}(\tau)$ is the velocity vector of the curve at parameter $\tau$. This length functional measures the total distance traveled along the curve, accounting for the anisotropic nature of our metric.
The energy of a curve, which is often more convenient for computational purposes, is defined as:
[…]
This energy functional penalizes curves that move in directions normal to the data manifold (when $\mathbf{s}(\gamma(\tau))^{T}\dot{\gamma}(\tau)\neq 0$) while imposing no additional cost for movement along tangential directions. We choose to optimize the energy functional rather than length because it eliminates the square root operation, making numerical optimization more stable.
Geodesics
Geodesics are curves with minimal length connecting two points in a Riemannian manifold. In our score-based metric space, geodesics are fundamental as they provide the optimal paths between points while naturally respecting the underlying data manifold structure.
Formally, a geodesic in the ambient space $\mathbb{R}^{N}$ is a curve with minimal length between two points $\mathbf{x}_{A}$ and $\mathbf{x}_{B}\in\mathbb{R}^{N}$. To find the shortest path between these points, we solve for the curve that minimizes the length or, equivalently and more conveniently, the energy functional
[…]
Where $\gamma^{*}:=\gamma^{*}(\tau)$ represents the optimal geodesic curve connecting $\mathbf{x}_{A}$ and $\mathbf{x}_{B}$.
Geodesic Computation Algorithm
While the geodesic equation provides a mathematical foundation, directly solving this minimization problem in high-dimensional spaces is challenging due to the computational complexity and the irregular behavior of score functions in data-sparse regions. Our approach leverages diffusion models to address these challenges through a three-stage process.
First, we apply controlled noise perturbation by sampling $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ and computing $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ for both endpoints. This forward diffusion process serves two purposes: it smooths the optimization landscape by regularizing the energy functional and ensures that identical noise is applied to both endpoints, preserving their relative positioning while making the optimization more stable. The noise level $t$ is a hyperparameter that controls the smoothness of the optimization landscape—higher values provide more stability but potentially less accurate manifold adherence.
With the score function $\mathbf{s}(\mathbf{x})$ obtained from the diffusion model at timestep $t$, we construct our metric tensor (Eq. 4) using a fixed scale parameter $\lambda$. The optimal choice of $\lambda$ depends on both the dataset characteristics and the specific noise level $t$ (see Section 4). Next, we discretize the energy functional and solve it numerically (for more details see B). We represent the path as a sequence of points $\gamma=\{\gamma_{0},\gamma_{1},...,\gamma_{n}\}$, with $\gamma_{0}=\mathbf{x}_{A}$ and $\gamma_{n}=\mathbf{x}_{B}$. The discrete energy becomes:
$\mathcal{E}[\gamma]\approx\frac{1}{2}\sum_{i=0}^{n-1}\|(\gamma_{i+1}-\gamma_{i})\|^{2}_{g(\gamma_{i})}=\frac{1}{2}\sum_{i=0}^{n-1}\left[\|(\gamma_{i+1}-\gamma_{i})\|^{2}+\lambda(\mathbf{s}(\gamma_{i})^{T}(\gamma_{i+1}-\gamma_{i}))^{2}\right]$ (8)
We minimize this energy using Riemannian gradient descent methods [24] that respect the curved geometry defined by our metric (for more details see Appendix B and Algorithm 2). The gradient computation takes into account how changes in each interior point affect both its contribution to the energy and the score-based metric at that point. We initialize the path with linear interpolation between the endpoints and iteratively refine it until convergence or a maximum iteration count is reached (see Appendix for experiment specific details). Finally, we map the optimized path in noise space back to image space using the denoising capabilities of the diffusion model. Each point along the geodesic is denoised from timestep $t$ back to $t^{\prime}=0$.
The three-stage process—noise addition, geodesic optimization, and denoising—creates paths that naturally follow the data manifold while remaining computationally tractable for high-dimensional image data.
Manifold-Aware Interpolation
Armed with our score-based metric tensor and geodesic computation algorithm, we can now perform interpolation between data points that respects the underlying manifold structure. Unlike conventional approaches such as linear interpolation (LERP) [10], spherical interpolation (SLERP) [25, 26, 27] or Noise Diffusion [28] that often cut through low-density regions of the data space, our geodesic-based interpolation follows the natural contours of the data manifold. This quality improvement (see Section 4) stems from the path’s adherence to the manifold structure, which prevents interpolation through low-density "off-manifold" regions where the diffusion model has not been trained.
Manifold-Aware Extrapolation
While interpolation connects two known endpoints, extrapolation extends a path beyond the observed data, requiring a different approach that respects the manifold without a target endpoint. We implement manifold-aware extrapolation through guided walking by the score-metric tensor with momentum.
Given a geodesic path ending at point $\mathbf{x}_{B}$, we extrapolate by iteratively computing new points:
$\mathbf{x}_{i+1}=\mathbf{x}_{i}+\mathbf{d}_{i}$ (9)
where $\mathbf{d}_{i}$ is a direction vector computed as a weighted combination of three components:
$\mathbf{d}_{i}=(1-\varepsilon)\cdot\mathbf{m}_{i}+\varepsilon\cdot\mathbf{s}(\mathbf{x}_{i})$ (10)
Here $\mathbf{m}_{i}$ is the momentum term that maintains trajectory consistency, $\mathbf{s}(\mathbf{x}_{i})$ is the score function guiding the path toward the data manifold. The $\varepsilon$ parameter controls the influence of manifold guidance. The momentum vector is updated using an exponential moving average:
$\mathbf{m}_{i+1}=\beta\cdot\mathbf{m}_{i}+(1-\beta)\cdot\mathbf{d}_{i}$ (11)
where $\beta$ controls how strongly the extrapolation maintains its previous direction. We initialize $\mathbf{m}_{0}$ from the tangent direction at the endpoint of the geodesic path (for more details see E). This approach offers several advantages: it maintains coherent progression by preserving directional momentum; it adheres to the manifold through score guidance. With computational complexity of $\mathcal{O}(n\cdot d)$ per step and requiring only one score evaluation per iteration.
4 Results
4.1 Embedded sphere
To validate our approach, we first consider a 2-sphere embedded in $\mathbb{R}^{100}$. We generate samples from a von Mises-Fisher distribution on $\mathbb{S}^{2}\subset\mathbb{R}^{3}$ (Figure 2A) and construct an isometric embedding – through QR decomposition – into $\mathbb{R}^{100}$ , with points representable as 10×10 pixel images (Figure 2B). This controlled testbed offers analytical tractability while mimicking high-dimensional image data.
We train a diffusion model on the embedded samples and extract the score function $\mathbf{s}(\mathbf{x})$. As shown in Figure 2C, the learned score vectors align with normal vectors of the sphere (quantitative validation in Supplementary Material), replicating the theoretical prediction that score functions are normal to the data manifold [13]. Using these score functions, we compute geodesics and compare with linear interpolation. Figure 2D-F shows that while linear paths cut through the sphere (producing blurred intermediate images), our geodesics follow the manifold’s surface (maintaining high sample quality throughout). Similarly, for extrapolation, our approach follows the sphere’s curvature and preserves image structure, while linear extrapolation quickly departs from the manifold, causing severe distortions.
Figure 2: Embedded sphere experiments: (A) von Mises-Fisher distribution on a 2-sphere. (B) Example 10×10 pixel images from the embedding. (C) Score vectors (purple) point normally to the manifold. (D) Linear interpolation (orange) versus geodesics (blue) between endpoints. (E) Extrapolation comparison with geodesics (purple) following the manifold versus linear paths (green) departing from it. (F) Image space results showing our geodesic paths maintain sample quality while linear paths produce blurring (interpolation) or artifacts (extrapolation).
4.2 Rotated MNIST
To evaluate whether our score-based metric captures perceptually meaningful transformations, we conducted experiments on a modified version of the MNIST dataset where the transformation structure is well-defined. We created a custom Rotated MNIST dataset by systematically rotating each digit by angles ranging from $0^{\circ}$ to $350^{\circ}$ in $10^{\circ}$ increments, producing 36 rotated versions of each original digit. We then trained a denoising diffusion model on this dataset using a UNet-2D [29] architecture with attention blocks. The model was trained for 100 epochs with a batch size of 256 (see more details in Appendix). Our hypothesis is that this trained model implicitly learns the data manifold, including the rotational structure of the dataset.
For interpolation experiments, we selected test set digit pairs with varying orientations, typically choosing the same digit rendered at different angles. Figure 3A compares our geodesic interpolation method against LERP, SLERP and Noise Diffusion. The results show that our geodesic approach produces trajectories that follow the rotational structure of the data manifold, creating smooth transformations that preserve digit identity while naturally rotating from one orientation to another. While LERP often produces unrealistic blending artifacts where intermediate frames superimpose features from both orientations rather than rotating continuously, our method follows the natural rotation transformation. Quantitatively, we measure the quality of interpolated frames using both PSNR and SSIM [30] metrics (Table 1), with our geodesic approach outperforming other methods in intermediate frame quality.
Table 1: Quality metrics on 100 Rotated MNIST test samples.
Metric LERP SLERP NoiseDiff Geodesic (Ours)
PSNR $\uparrow$ 14.08 $\pm$ 1.27 13.64 $\pm$ 1.33 13.67 $\pm$ 0.83 14.98 $\pm$ 1.29
SSIM $\uparrow$ 0.578 $\pm$ 0.062 0.572 $\pm$ 0.074 0.568 $\pm$ 0.054 0.650 $\pm$ 0.056
The extrapolation experiments, shown in Figure 3B, demonstrate the power of our manifold-guided approach for extending transformations beyond observed data points. Starting with two orientations of the same digit, we compute the geodesic path between them and then continue the transformation using our extrapolation algorithm. The results show that our method successfully continues the rotational trajectory, generating novel orientations that maintain digit identity and structure despite never having been explicitly observed in this sequence during training.
Figure 3: Rotated MNIST. A Interpolation Example (Best LERP by PSNR) comparing LERP, SLERP, Noise Diffusion and Geodesic (our method). B Three examples with our extrapolation method.
4.3 Stable Diffusion & MorphBench
To demonstrate our method’s applicability to state-of-the-art diffusion models, we conducted experiments with Stable Diffusion 2.1 [31], a powerful latent diffusion model trained on billions of image-text pairs. This allowed us to scale our approach on a more complex, higher-resolution image generation task than our previous experiments. Stable Diffusion operates in a compressed latent space rather than pixel space, making it an interesting test case for our score-based metric. The diffusion process occurs in a 4×64×64 latent space (representing 512×512 pixel, RGB images), where the data manifold has complex geometry reflecting natural image statistics. We apply our methodology by computing geodesics in this latent space using score estimates from the model’s denoising U-Net, then mapping the results back to image space using the model’s VAE decoder.
For evaluation, we utilize MorphBench [32], a comprehensive benchmark for assessing image morphing capabilities. MorphBench consists of 90 diverse image pairs organized into two categories: (1) metamorphosis between different objects (66 pairs) and (2) animation of the same object with different attributes (24 pairs). This diversity allows us to evaluate our method across varying transformation complexities.
Figure 4 (see Appendix G, for more examples) presents qualitative comparisons between our geodesic method and baseline approaches (LERP, SLERP, and Noise Diffusion). Our method generates smoother, more natural-looking transitions that better preserve image coherence during the morphing process. While LERP and SLERP often produce blurry intermediate frames with unrealistic composites of both source and target images, and Noise Diffusion frequently generates images with artifacts, our geodesic interpolation creates a more natural progression by following meaningful paths on the data manifold.
Table 2 provides a comprehensive quantitative evaluation. Our geodesic method achieves the best performance on LPIPS (0.3582 vs. 0.3613 for LERP), a learned perceptual similarity metric that correlates strongly with human judgments of visual quality [33]. We also outperform all baselines on distribution-level metrics, with the lowest FID (140.61 vs. 148.19 for LERP) [34] and KID scores (0.0863 vs. 0.0935 for LERP) [35]. These results indicate that our interpolated images are both perceptually coherent and maintain high sample quality that better matches the "true" data distribution. While LERP achieves slightly better results on pixel-level metrics (PSNR: 21.28 vs. 20.88; SSIM: 0.6274 vs. 0.6180), it’s well-established that these metrics often fail to capture perceptual quality [33], especially for high-resolution natural images.
Metric LERP SLERP NOISEDIFF GEODESIC (Ours)
avg_SSIM $\uparrow$ 0.627 $\pm$ 0.124 0.575 $\pm$ 0.144 0.404 $\pm$ 0.119 0.618 $\pm$ 0.117
avg_PSNR $\uparrow$ 21.28 $\pm$ 1.81 20.36 $\pm$ 2.25 16.73 $\pm$ 2.03 20.88 $\pm$ 1.78
avg_LPIPS $\downarrow$ 0.361 $\pm$ 0.077 0.396 $\pm$ 0.081 0.480 $\pm$ 0.046 0.358 $\pm$ 0.070
FID $\downarrow$ 148.2 $\pm$ 50.3 171.5 $\pm$ 53.6 271.8 $\pm$ 47.6 140.6 $\pm$ 51.3
KID $\downarrow$ 0.094 $\pm$ 0.052 0.110 $\pm$ 0.057 0.186 $\pm$ 0.064 0.086 $\pm$ 0.055
Table 2: Aggregated metrics for MorphBench across 90 image pairs.
Figure 4: MorphBench interpolation example with Stable Diffusion 2.1. Comparing LERP, SLERP, Noise Diffusion and Geodesic (Our method)
5 Discussion
In this work, we introduced a score-based Riemannian metric derived from diffusion models that captures the intrinsic geometry of the data manifold without requiring explicit parameterization. Our approach leverages the Stein score function to define a metric that naturally adapts to the manifold structure, enabling geometric computations that respect the underlying data distribution. Through experiments on synthetic data, MNIST, and complex natural images, we demonstrated that geodesics in our metric space correspond to perceptually meaningful paths between images, outperforming conventional interpolation methods.
While we used interpolation and extrapolation as validation applications, our approach provides a general framework for understanding and exploring the geometric structure learned by diffusion models. The strong performance on these tasks—despite not being specifically optimized for image morphing—suggests that our score-based metric naturally aligns with human perception of image similarity and transformation paths. This is evidenced by the consistent superiority of our geodesic interpolation approach on perceptual metrics like LPIPS and distribution metrics like FID and KID, indicating that our method produces images that not only look better individually but also follow more natural trajectories through the data space.
An interesting pattern emerged when comparing performance across different datasets. On Rotated MNIST, our method outperformed baselines across pixel-level metrics (PSNR and SSIM). However, on the higher-resolution MorphBench dataset with Stable Diffusion, our method dominated in perceptual metrics (LPIPS) and distribution metrics (FID, KID) but was slightly outperformed by LERP on pixel-level metrics. This difference can be attributed to our geodesic computations for Stable Diffusion occurring in the compressed latent space rather than the pixel space. The VAE decoder introduces additional variability when mapping back to pixels, which may affect pixel-perfect reconstruction while still preserving—and even enhancing—perceptual quality. This observation aligns with the established understanding that pixel-level metrics often fail to capture perceptual similarity in high-resolution natural images, instead favoring blurry but pixel-wise accurate reconstructions over sharper, perceptually preferable images [33].
Limitations. Despite its strengths, our approach introduces additional computational complexity compared to direct methods like LERP or SLERP. Our implementation typically requires several hundred optimization steps to converge, whereas alternative methods can be computed without any iterative optimization procedure.
Another limitation relates to the validation of extrapolation results. Although our experiments demonstrate that the extrapolated images maintain high quality and follow natural extensions of the geodesic paths, there are no established quantitative metrics to evaluate the quality of extrapolations in our settings, making objective comparisons challenging. Nevertheless, we view extrapolation as a data-driven exploration of the learned representation—a tool for discovering the implicit structure captured by diffusion models rather than a task requiring ground truth validation.
Future Work. One exciting direction would be applying our approach to semantic image editing tasks [36, 37]. By computing geodesic paths between edited and original images, our method could enable more natural and realistic transitions when performing attribute manipulations, style transfers, or object insertions.
Additionally, the computational efficiency of geodesic calculation could be improved through techniques like neural surrogate models [38] that directly predict geodesic paths. Such models could enable real-time interactive editing tools that allow artists and designers to explore the learned manifold of a diffusion model while maintaining high sample quality throughout the editing process.
Finally, our score-based metric provides a new lens for analyzing the geometry learned by diffusion models. Future work could explore using this geometric perspective to better understand model behavior, detect biases in learned distributions, or visualize how the data manifold evolves during training. By continuing to develop the connections between diffusion models and Riemannian geometry, we can gain deeper insights into how these powerful generative models represent complex data distributions.
Appendix A Proof of Riemannian Metric Properties
To prove that $g(\boldsymbol{x})=\mathbf{I}+\lambda\cdot\boldsymbol{s}(\boldsymbol{x})\boldsymbol{s}(\boldsymbol{x})^{T}$ defines a valid Riemannian metric, we need to show that it satisfies the two fundamental properties:
Symmetry: For any point $\boldsymbol{x}$, $g(\boldsymbol{x})$ must be symmetric.
$g(\boldsymbol{x})^{T}$ $=(\mathbf{I}+\lambda\cdot\boldsymbol{s}(\boldsymbol{x})\boldsymbol{s}(\boldsymbol{x})^{T})^{T}$
$=\mathbf{I}^{T}+\lambda\cdot(\boldsymbol{s}(\boldsymbol{x})\boldsymbol{s}(\boldsymbol{x})^{T})^{T}$
$=\mathbf{I}+\lambda\cdot(\boldsymbol{s}(\boldsymbol{x})^{T})^{T}\boldsymbol{s}(\boldsymbol{x})^{T}$
$=\mathbf{I}+\lambda\cdot\boldsymbol{s}(\boldsymbol{x})\boldsymbol{s}(\boldsymbol{x})^{T}$
$=g(\boldsymbol{x})$
Positive-definiteness: For any point $\boldsymbol{x}$ and any non-zero vector $\boldsymbol{v}$, we must have $\boldsymbol{v}^{T}g(\boldsymbol{x})\boldsymbol{v}>0$.
$\boldsymbol{v}^{T}g(\boldsymbol{x})\boldsymbol{v}$ $=\boldsymbol{v}^{T}(\mathbf{I}+\lambda\cdot\boldsymbol{s}(\boldsymbol{x})\boldsymbol{s}(\boldsymbol{x})^{T})\boldsymbol{v}$
$=\boldsymbol{v}^{T}\mathbf{I}\boldsymbol{v}+\lambda\cdot\boldsymbol{v}^{T}\boldsymbol{s}(\boldsymbol{x})\boldsymbol{s}(\boldsymbol{x})^{T}\boldsymbol{v}$
$=\|\boldsymbol{v}\|^{2}+\lambda\cdot(\boldsymbol{s}(\boldsymbol{x})^{T}\boldsymbol{v})^{2}$
Since $\lambda>0$ (by construction) and $\|\boldsymbol{v}\|^{2}>0$ for any non-zero vector $\boldsymbol{v}$, we have $\boldsymbol{v}^{T}g(\boldsymbol{x})\boldsymbol{v}>0$. The second term $\lambda\cdot(\boldsymbol{s}(\boldsymbol{x})^{T}\boldsymbol{v})^{2}$ is always non-negative, further ensuring positive-definiteness.
Thus, the Stein score metric tensor $g(\boldsymbol{x})=\mathbf{I}+\lambda\cdot\boldsymbol{s}(\boldsymbol{x})\boldsymbol{s}(\boldsymbol{x})^{T}$ satisfies all the requirements of a Riemannian metric on $\mathbb{R}^{N}$.
Appendix B Geodesic Algorithm Details
B.1 Discretization and Numerical Integration
To compute geodesics numerically, we discretize the continuous energy functional using the midpoint rule for integration. This approach provides second-order accuracy and better stability compared to first-order methods such as Euler integration.
Given endpoints $\boldsymbol{x}_{A}$ and $\boldsymbol{x}_{B}$, we represent the geodesic path as a sequence of $n+1$ points:
$\gamma=\{\gamma_{0},\gamma_{1},\ldots,\gamma_{n}\}$ (12)
where $\gamma_{0}=\boldsymbol{x}_{A}$ and $\gamma_{n}=\boldsymbol{x}_{B}$ are fixed, and the interior points $\{\gamma_{1},\ldots,\gamma_{n-1}\}$ are optimized.
For each segment of the discretized path, we:
  1. 1.
    Compute the velocity vector: $\boldsymbol{v}_{i}=\gamma_{i+1}-\gamma_{i}$
  2. 2.
    Calculate the midpoint: $\boldsymbol{m}_{i}=\frac{1}{2}(\gamma_{i+1}+\gamma_{i})$
  3. 3.
    Evaluate the score function at the midpoint: $\boldsymbol{s}_{i}=\boldsymbol{s}(\boldsymbol{m}_{i})$
  4. 4.
    Compute the segment’s contribution to the energy functional
The discretized energy functional under the midpoint approximation becomes:
$\mathcal{E}[\gamma]\approx\frac{1}{2}\sum_{i=0}^{n-1}\|\boldsymbol{v}_{i}\|^{2}_{g(\boldsymbol{m}_{i})}=\frac{1}{2}\sum_{i=0}^{n-1}\left[\|\boldsymbol{v}_{i}\|^{2}+\lambda(\boldsymbol{s}_{i}^{T}\boldsymbol{v}_{i})^{2}\right]$ (13)
To improve the optimization stability, we incorporate additional regularization terms:
$\mathcal{E}_{\text{total}}[\gamma]=\mathcal{E}[\gamma]+\lambda_{\text{smooth}}\mathcal{R}_{\text{smooth}}[\gamma]+\lambda_{\text{mono}}\mathcal{R}_{\text{mono}}[\gamma]$ (14)
where:
$\mathcal{R}_{\text{smooth}}[\gamma]$ $=\sum_{i=1}^{n-1}\|\gamma_{i+1}-2\gamma_{i}+\gamma_{i-1}\|^{2}$ (15)
$\mathcal{R}_{\text{mono}}[\gamma]$ $=\sum_{i=1}^{n-1}\max(0,\|\gamma_{i}-\gamma_{n}\|-\|\gamma_{i-1}-\gamma_{n}\|)$ (16)
The smoothness term penalizes high curvature in the path, while the monotonicity term encourages the path to make consistent progress toward the endpoint.
B.2 Riemannian Adam Optimization
To minimize the discretized energy functional, we use a Riemannian extension of the Adam optimizer that properly accounts for the curved geometry defined by our metric. The algorithm proceeds as follows:
Algorithm 1 Riemannian Adam for Geodesic Optimization
1:Input: Initial path $\gamma=\{\gamma_{0},\gamma_{1},\ldots,\gamma_{n}\}$ with fixed endpoints
2:Hyperparameters: Learning rate $\alpha$, momentum parameters $\beta_{1},\beta_{2}$, stability parameter $\epsilon$
3:Initialize: Moment vectors $\boldsymbol{m}_{i}=\boldsymbol{0}$, $\boldsymbol{v}_{i}=\boldsymbol{0}$ for $i\in\{1,\ldots,n-1\}$
4:$t\leftarrow 0$
5:while not converged do
6:     $t\leftarrow t+1$
7:     Compute $\mathcal{E}_{\text{total}}[\gamma]$ and its gradients $\{\nabla_{\gamma_{i}}\mathcal{E}_{\text{total}}\}_{i=1}^{n-1}$
8:     for $i=1$ to $n-1$ do
9:         $\boldsymbol{g}_{i}\leftarrow\nabla_{\gamma_{i}}\mathcal{E}_{\text{total}}$
10:         $\widetilde{\boldsymbol{g}}_{i}\leftarrow$ RiemannianGradient$(\boldsymbol{g}_{i},\gamma_{i},g)$triangleright Convert Euclidean to Riemannian gradient
11:         $\boldsymbol{m}_{i}\leftarrow\beta_{1}\boldsymbol{m}_{i}+(1-\beta_{1})\widetilde{\boldsymbol{g}}_{i}$
12:         $\boldsymbol{v}_{i}\leftarrow\beta_{2}\boldsymbol{v}_{i}+(1-\beta_{2})\widetilde{\boldsymbol{g}}_{i}^{2}$
13:         $\hat{\boldsymbol{m}}_{i}\leftarrow\boldsymbol{m}_{i}/(1-\beta_{1}^{t})$triangleright Bias correction
14:         $\hat{\boldsymbol{v}}_{i}\leftarrow\boldsymbol{v}_{i}/(1-\beta_{2}^{t})$
15:         $\boldsymbol{\eta}_{i}\leftarrow\alpha\cdot\hat{\boldsymbol{m}}_{i}/(\sqrt{\hat{\boldsymbol{v}}_{i}}+\epsilon)$triangleright Update direction
16:         $\gamma_{i}^{\text{old}}\leftarrow\gamma_{i}$
17:         $\gamma_{i}\leftarrow\gamma_{i}-\boldsymbol{\eta}_{i}$triangleright Update position
18:         $\boldsymbol{m}_{i}\leftarrow$ ParallelTransport$(\boldsymbol{m}_{i},\gamma_{i}^{\text{old}},\gamma_{i},g)$triangleright Transport momentum vector
19:     end for
20:     if convergence criteria met then
21:         break
22:     end if
23:end while
24:Return optimized path $\gamma$
The key components that extend Adam to Riemannian manifolds are:
Riemannian Gradient: The Euclidean gradient $\boldsymbol{g}$ is converted to a Riemannian gradient $\widetilde{\boldsymbol{g}}$ using:
$\widetilde{\boldsymbol{g}}=g(\boldsymbol{x})^{-1}\boldsymbol{g}$ (17)
For our metric tensor $g(\boldsymbol{x})=\mathbf{I}+\lambda\cdot\boldsymbol{s}(\boldsymbol{x})\boldsymbol{s}(\boldsymbol{x})^{T}$, we can efficiently compute this using the Sherman-Morrison formula:
$\widetilde{\boldsymbol{g}}=\boldsymbol{g}-\frac{\lambda\cdot(\boldsymbol{s}^{T}\boldsymbol{g})\cdot\boldsymbol{s}}{1+\lambda\cdot(\boldsymbol{s}^{T}\boldsymbol{s})}$ (18)
Parallel Transport: To properly preserve momentum information when moving between points on the manifold, we parallel transport the momentum vectors along the update direction:
[…]
This first-order approximation of parallel transport is sufficient for our purposes and maintains the directional information of the momentum vectors as they move along the curved manifold.
Through this optimization process, we obtain a discretized geodesic path that minimizes the energy functional while respecting the Riemannian geometry induced by our score-based metric tensor.
Appendix C Geodesic Computation Algorithm
We complement the geodesic algorithm description in the main paper with Algorithm 2
Algorithm 2 Geodesic Computation with Score-Based Riemannian Metric
1:Input: Points $\boldsymbol{p}$, $\boldsymbol{q}$, diffusion timestep $t$, number of segments $n$, scale parameter $\lambda$
2:Output: Geodesic path $\gamma=\{\gamma_{0},\gamma_{1},\ldots,\gamma_{n}\}$ with $\gamma_{0}=\boldsymbol{p}$, $\gamma_{n}=\boldsymbol{q}$
3:Forward diffusion stage:
4:Sample noise $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$
5:$\tilde{\boldsymbol{p}}\leftarrow\sqrt{\alpha_{t}}\boldsymbol{p}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}$
6:$\tilde{\boldsymbol{q}}\leftarrow\sqrt{\alpha_{t}}\boldsymbol{q}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}$
7:Initialize path:
8:$\gamma_{i}\leftarrow(1-\lambda_{i})\tilde{\boldsymbol{p}}+\lambda_{i}\tilde{\boldsymbol{q}}$ for $i\in\{0,1,\ldots,n\}$, $\lambda_{i}=i/n$
9:Interior points $\Gamma\leftarrow\{\gamma_{1},\gamma_{2},\ldots,\gamma_{n-1}\}$
10:Define score-based metric tensor:
11:$g_{\boldsymbol{x}}(\boldsymbol{u},\boldsymbol{v})=\boldsymbol{u}^{T}\boldsymbol{v}+\lambda\cdot(\boldsymbol{s}(\boldsymbol{x})^{T}\boldsymbol{u})(\boldsymbol{s}(\boldsymbol{x})^{T}\boldsymbol{v})$
12:Optimize path:
13:Initialize RiemannianAdam optimizer with interior points $\Gamma$
14:for iteration $=1$ to max_iterations do
15:     // Compute energy using midpoint discretization
16:     $\mathcal{E}\leftarrow 0$
17:     for $i=0$ to $n-1$ do
18:         $\boldsymbol{v}_{i}\leftarrow\gamma_{i+1}-\gamma_{i}$triangleright Segment velocity
19:         $\boldsymbol{m}_{i}\leftarrow\frac{1}{2}(\gamma_{i+1}+\gamma_{i})$triangleright Segment midpoint
20:         $\mathcal{E}\leftarrow\mathcal{E}+\frac{1}{2}\|\boldsymbol{v}_{i}\|^{2}+\frac{\lambda}{2}(\boldsymbol{s}(\boldsymbol{m}_{i})^{T}\boldsymbol{v}_{i})^{2}$triangleright Energy contribution
21:     end for
22:     // Add regularization terms
23:     $\mathcal{E}_{\text{smooth}}\leftarrow\lambda_{\text{smooth}}\sum_{i=1}^{n-1}\|\gamma_{i+1}-2\gamma_{i}+\gamma_{i-1}\|^{2}$triangleright Smoothness
24:     $\mathcal{E}_{\text{mono}}\leftarrow\lambda_{\text{mono}}\sum_{i=1}^{n-1}\text{ReLU}(\|\gamma_{i}-\gamma_{n}\|-\|\gamma_{i-1}-\gamma_{n}\|)$triangleright Monotonicity
25:     $\mathcal{E}_{\text{total}}\leftarrow\mathcal{E}+\mathcal{E}_{\text{smooth}}+\mathcal{E}_{\text{mono}}$
26:     Compute gradients of $\mathcal{E}_{\text{total}}$ with respect to interior points $\Gamma$
27:     Update interior points using Riemannian gradient step
28:     if convergence criteria met then
29:         break
30:     end if
31:end for
32:Reverse diffusion:
33:for $i=0$ to $n$ do
34:     if $i=0$ or $i=n$ then
35:         $\hat{\gamma}_{i}\leftarrow$ original clean point ($\boldsymbol{p}$ or $\boldsymbol{q}$)
36:     else
37:         $\hat{\gamma}_{i}\leftarrow$ Denoise($\gamma_{i}$, from $t$ to 00)
38:     end if
39:end for
40:return $\hat{\gamma}=\{\hat{\gamma}_{0},\hat{\gamma}_{1},\ldots,\hat{\gamma}_{n}\}$
Appendix D Interpolation
We complement the interpolation algorithm description in the main paper with Algorithm 3
Algorithm 3 Manifold-Aware Interpolation
1:Input: Points $\boldsymbol{p}$, $\boldsymbol{q}$, number of interpolation points $n$, timestep $t$
2:Output: Interpolated sequence $\{\boldsymbol{y}_{0},\boldsymbol{y}_{1},\ldots,\boldsymbol{y}_{n}\}$
3:Stage 1: Forward Diffusion
4:Sample noise $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$triangleright Same noise for consistency
5:$\tilde{\boldsymbol{p}}\leftarrow\sqrt{\alpha_{t}}\boldsymbol{p}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}$
6:$\tilde{\boldsymbol{q}}\leftarrow\sqrt{\alpha_{t}}\boldsymbol{q}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}$
7:Stage 2: Geodesic Computation in Noise Space
8:Set up score extractor $s_{\theta}$ at timestep $t$
9:Set up Stein metric tensor with adaptive scaling
10:Compute reference scores at endpoints
11:Initialize path via linear interpolation
12:Optimize path using Algorithm 2 (Geodesic Computation)
13:Stage 3: Backward Diffusion (Denoising)
14:for $i=0$ to $n$ do
15:     if $i=0$ then
16:         $\boldsymbol{y}_{i}\leftarrow\boldsymbol{p}$triangleright Use original clean endpoint
17:     else if $i=n$ then
18:         $\boldsymbol{y}_{i}\leftarrow\boldsymbol{q}$triangleright Use original clean endpoint
19:     else
20:         $\boldsymbol{y}_{i}\leftarrow$ DenoisingDiffusion($\tilde{\gamma}_{i}$, from $t$ to 00)
21:     end if
22:end for
23:return $\{\boldsymbol{y}_{0},\boldsymbol{y}_{1},\ldots,\boldsymbol{y}_{n}\}$
Appendix E Extrapolation
We complement the interpolation algorithm description in the main paper with Algorithm 4
Algorithm 4 Single-Direction Extrapolation
1:Initialize $\boldsymbol{x}_{\text{current}}\leftarrow\boldsymbol{q}$
2:Compute initial direction from path segments:
3:$\boldsymbol{m}\leftarrow\sum_{i=1}^{\min(3,n/4)}w_{i}\cdot(\gamma_{n-i+1}-\gamma_{n-i})/\sum_{i}w_{i}$
[…]
5:for $i=1$ to num_steps do
6:     Compute score at current point: $\boldsymbol{s}\leftarrow\mathbf{s}(\boldsymbol{x}_{\text{current}})$
7:     Compute direction: $\boldsymbol{d}\leftarrow(1-\varepsilon)\cdot\boldsymbol{m}+\varepsilon\cdot\boldsymbol{s}$
[…]
9:     Update position: $\boldsymbol{x}_{\text{next}}\leftarrow\boldsymbol{x}_{\text{current}}+\boldsymbol{d}$
10:     Update momentum: $\boldsymbol{m}\leftarrow\beta\cdot\boldsymbol{m}+(1-\beta)\cdot\boldsymbol{d}$
11:     $\boldsymbol{x}_{\text{current}}\leftarrow\boldsymbol{x}_{\text{next}}$
12:     Add $\boldsymbol{x}_{\text{current}}$ to extrapolation path
13:end for
Appendix F Rotated MNIST
After setting up the dataset as in Section 4.2, we trained a DDPM [10] Model in PyTorch (version 2.7 and cuda 12.6) by exploiting the Diffusers library [39] and Accellerate library [40] to parallelize training on 7 NVIDIA A100 GPUs.
U-Net Network Architecture
We employed a U-Net model with the following configuration:
  • Input/Output: The model accepts 32$\times$32 grayscale images (single channel) and predicts noise with matching dimensions.
  • Depth: Four resolution levels with downsampling and upsampling operations.
  • Channel Dimensions: Channel counts of (64, 128, 256, 256) at each respective resolution level.
  • Block Structure: Each resolution level contains 2 ResNet blocks.
  • Attention Mechanism: Self-attention blocks at the third level of both encoder and decoder paths to capture global relationships, which is crucial for understanding rotational structure.
  • Downsampling Path: (DownBlock2D, DownBlock2D, AttnDownBlock2D, DownBlock2D)
  • Upsampling Path: (UpBlock2D, AttnUpBlock2D, UpBlock2D, UpBlock2D)
The total parameter count of the model is approximately 30 million parameters.
The training procedure was as follows:
  • Optimizer: AdamW with learning rate $1\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$
  • Learning rate scheduler: Cosine schedule with warmup (500 steps)
  • Batch size: 256
  • Training duration: 100 epochs
  • Loss function: Mean squared error (MSE) between predicted and true noise
  • Precision: Mixed precision training with bfloat16
  • Training Time: 12 hours
Inference and geodesic optimization was then done on single GPU. For Table 1 we randomly sampled 100 digits from the MNIST test set (not used for training the diffusion models), fixed $t=400$ diffusion process steps, and run LERP, SLERP, NoiseDiffusion and our method (Geodesic); after that we denoised back to image space and computed PSNR and SSIM.
The geodesic optimization - with 8 points, as in Figure 3 A - took 2 minutes on average per iteration (2000 maximum amount of optimization epochs with 250 patience).
After that we selected a handful of test samples, computed the geodesic interpolation between a reference base angle and this angle + 30°. We then extrapolated for N = 5 steps (results are shown in Figure 3 B). Image extrapolation was extremely fast ($\simeq 1$ second for 5 steps).
Appendix G Stable Diffusion & MorphBench
We employed the pre-trained Stable Diffusion 2.1. implementation available in the Diffusers library [39], running with PyTorch 2.7 and cuda 12.6 on a NVIDIA A100 GPU (40GB VRAM).
Geodesic computation with 10 points, 2000 iterations (patience 250 epochs), tipically required 20 mins per image interpolation. We employed t = 400 (diffusion steps)
For our experiments with Stable Diffusion 2.1, we used unconditional generation by providing an empty text prompt (""), allowing the model to focus solely on the image manifold structure without text-based guidance.
MorphBench Performances
Examining the method performance across dataset categories reveals interesting patterns. For the animation subset (Table 3 in Appendix), our method excels in perceptual metrics (best LPIPS at 0.3402 and KID at 0.0901), while LERP performs marginally better in pixel metrics and FID. For the more challenging metamorphosis transformations (Table 4 in Appendix), our geodesic approach demonstrates clear advantages in distribution-level metrics (best FID at 145.94 and KID at 0.0848). This superiority in metamorphosis scenarios highlights our method’s effectiveness for complex transformations between different objects, where properly navigating the intrinsic manifold geometry is most critical.
Additionally we report more examples of interpolation comparison, see Figure 5,  6,  7
Metric LERP SLERP NOISEDIFFUSION GEODESIC
avg_ssim $\uparrow$ 0.6251 $\pm$ 0.1084 $0$$\pm$ 0.1198 0.4202 $\pm$ 0.1152 0.6261 $\pm$ 0.1026
avg_psnr $\uparrow$ 21.23 $\pm$ 2.08 20.72 $\pm$ 2.37 17.09 $\pm$ 2.31 20.98 $\pm$ 2.05
avg_lpips $\downarrow$ 0.3467 $\pm$ 0.0749 0.3652 $\pm$ 0.0816 0.4577 $\pm$ 0.0584 0.3402 $\pm$ 0.0679
fid $\downarrow$ 125.89 $\pm$ 44.13 142.08 $\pm$ 49.26 253.52 $\pm$ 54.07 127.29 $\pm$ 49.95
kid $\downarrow$ 0.0947 $\pm$ 0.0652 0.1059 $\pm$ 0.0716 0.1925 $\pm$ 0.0791 0.0901 $\pm$ 0.0659
Table 3: Aggregated metrics for Animation dataset across 24 image pairs.
Metric LERP SLERP NOISEDIFFUSION GEODESIC
avg_ssim $\uparrow$ 0.6282 $\pm$ 0.1290 0.5653 $\pm$ 0.1517 0.3969 $\pm$ 0.1201 0.6148 $\pm$ 0.1222
avg_psnr $\uparrow$ 21.29 $\pm$ 1.69 20.21 $\pm$ 2.19 16.58 $\pm$ 1.89 20.83 $\pm$ 1.65
avg_lpips $\downarrow$ 0.3672 $\pm$ 0.0774 0.4076 $\pm$ 0.0779 0.4891 $\pm$ 0.0362 0.3653 $\pm$ 0.0694
fid $\downarrow$ 157.11 $\pm$ 49.91 183.24 $\pm$ 50.73 279.13 $\pm$ 42.53 145.94 $\pm$ 50.85
kid $\downarrow$ 0.0930 $\pm$ 0.0457 0.1120 $\pm$ 0.0493 0.1836 $\pm$ 0.0560 0.0848 $\pm$ 0.0496
Table 4: Aggregated metrics for Metamorphosis dataset across 66 image pairs.
Figure 5: Stable Diffusion. Interpolation Example vs LERP, SLERP and Noise Diffusion.
Figure 6: Stable Diffusion. Interpolation Example vs LERP, SLERP and Noise Diffusion.
Figure 7: Stable Diffusion. Interpolation Example vs LERP, SLERP and Noise Diffusion.